# Empowering Source-Free Domain Adaptation via MLLM-Guided Reliability-Based Curriculum Learning

### v2

@@ 4.3 Multi-hot Masking Refinement (MMR)
5 Experiments
Datasets. We evaluate our method on three standard benchmark datasets: Office-Home, DomainNet-126, and VisDA-C 2017. Office-Home [32] has 4 domains – Real (R), Clipart (C), Art (A), and Product (P), encompassing 65 classes with a total of 15.5k images. VisDA [26] is a large-scale synthetic-to-real object recognition dataset, where the source domain includes 152k synthetic images and the target domain contains about 55k real object images across 12 classes. DomainNet [27] is a challenging large-scale, featuring 6 domains with a total of around 600k images across 345 classes. We follow the standard DomainNet-126 setup with 145k images from 126 classes, sampled from four domains, Clipart (C), Painting (P), Real (R), Sketch (S).
Model details. Following [41, 20, 39], we use ResNet-101 [8] for VisDA and ResNet-50 for Office-Home and DomainNet. For VisDA and Office-Home, we adopt pre-trained source models from SHOT [20], while for DomainNet, we train source models from scratch following [23]. Similar to DIFO [41], we also report results with ViT-B/32 backbone (RCL-ViT). For main results, we use the strongest open-source MLLMs for RCL: LLaVA-v1.6-34B [25], ShareGPT4V-13B [2], and InstructBLIP-T5-XXL [4].
Training Details. The parameters used in the training process of RCL are shown in Table 1. We use the Adam optimizer [14] for RCL training. All experiments were conducted using PyTorch on NVIDIA A100 GPUs.
Table 1: RCL Training Parameters per stage (RKT, SMKE, MMR).
Office-Home DomainNet VisDA
RKT SMKE MMR RKT SMKE MMR RKT SMKE MMR
learning rate 1e-04 1e-05 1e-05 1e-05 1e-05 1e-05 1e-05 1e-06 1e-06
$\tau$- 0.7 0.95 -- 0.7 0.9 -- 0.7 0.6
batch size 64 256 128 64 256 64 64 256 256
max iter 3000 5000 5000 8000 10000 5000 6000 6000 5000
Table 2: Accuracy (%) on Office-Home dataset. SF: source-free, CP, ViT: method uses CLIP, ViT. We highlight the best result and underline the second-best one. (*) represents pre-trained CLIP/MLLM zero-shot performance with proposed STS.
[…]
Source - 44.7 64.2 69.4 48.3 57.9 60.3 49.5 40.3 67.2 59.7 45.6 73.0 56.7
DAPL-RN [7] 54.1 84.3 84.8 74.4 83.7 85.0 74.5 54.6 84.8 75.2 54.7 83.8 74.5
PADCLIP-RN [15] 57.5 84.0 83.8 77.8 85.5 84.7 76.3 59.2 85.4 78.1 60.2 86.7 76.6
ADCLIP-RN [33] 55.4 85.2 85.6 76.1 85.8 86.2 76.7 56.1 85.4 76.8 56.1 85.5 75.9
SHOT [20] 56.7 77.9 80.6 68.0 78.0 79.4 67.9 54.5 82.3 74.2 58.6 84.5 71.9
NRC [45] 57.7 80.3 82.0 68.1 79.8 78.6 65.3 56.4 83.0 71.0 58.6 85.6 72.2
GKD [34] 56.5 78.2 81.8 68.7 78.9 79.1 67.6 54.8 82.6 74.4 58.5 84.8 72.2
AaD [47] 59.3 79.3 82.1 68.9 79.8 79.5 67.2 57.4 83.1 72.1 58.5 85.4 72.7
AdaCon [1] 47.2 75.1 75.5 60.7 73.3 73.2 60.2 45.2 76.6 65.6 48.3 79.1 65.0
CoWA [17] 56.9 78.4 81.0 69.1 80.0 79.9 67.7 57.2 82.4 72.8 60.5 84.5 72.5
SCLM [36] 58.2 80.3 81.5 69.3 79.0 80.7 69.0 56.8 82.7 74.7 60.6 85.0 73.0
ELR [48] 58.4 78.7 81.5 69.2 79.5 79.3 66.3 58.0 82.6 73.4 59.8 85.1 72.6
PLUE [23] 49.1 73.5 78.2 62.9 73.5 74.5 62.2 48.3 78.6 68.6 51.8 81.5 66.9
TPDS [38] 59.3 80.3 82.1 70.6 79.4 80.9 69.8 56.8 82.1 74.5 61.2 85.3 73.5
C-SFDA [13] 60.3 80.2 82.9 69.3 80.1 78.8 67.3 58.1 83.4 73.6 61.3 86.3 73.5
PSAT-GDA [39] 73.1 88.1 89.2 82.1 88.8 88.9 83.0 72.0 89.6 83.3 73.7 91.3 83.6
LCFD-C-RN [40] 60.1 85.6 86.2 77.2 86.0 86.3 76.6 61.0 86.5 77.5 61.4 86.2 77.6
LCFD-C-B32 [40] 72.3 89.8 89.9 81.1 90.3 89.5 80.1 71.5 89.8 81.8 72.7 90.4 83.3
DIFO-C-RN [41] 62.6 87.5 87.1 79.5 87.9 87.4 78.3 63.4 88.1 80.0 63.3 87.7 79.4
DIFO-C-B32 [41] 70.6 90.6 88.8 82.5 90.6 88.8 80.9 70.1 88.9 83.4 70.5 91.2 83.1
CLIP-RN [30]* - 51.7 85.0 83.7 69.3 85.0 83.7 69.3 51.7 83.7 69.3 51.7 85.0 72.4
LLaVA-34B [25]* (w/ STS) - 78.3 93.7 89.5 87.0 93.7 89.5 87.0 78.3 89.5 87.0 78.3 93.7 87.2
InstBLIP-XXL [4]* (w/ STS) - 82.0 91.6 88.8 82.2 91.6 88.8 82.2 82.0 88.8 82.2 82.0 91.6 86.2
ShrGPT4V-13B [2]* (w/ STS) - 66.7 85.8 84.8 83.2 85.8 84.8 83.2 66.7 84.8 83.2 66.7 85.8 80.1
Table 3: Accuracy (%) on DomainNet dataset.
[…]
Source - 42.6 53.7 51.9 52.9 66.7 51.6 49.1 56.8 43.9 60.9 48.6 53.2 52.7
DAPL-RN [7] 72.4 87.6 65.9 72.7 87.6 65.6 73.2 72.4 66.2 73.8 72.9 87.8 74.8
ADCLIP-RN [15] 71.7 88.1 66.0 73.2 86.9 65.2 73.6 73.0 68.4 72.3 74.2 89.3 75.2
SHOT [20] 63.5 78.2 59.5 67.9 81.3 61.7 67.7 67.6 57.8 70.2 64.0 78.0 68.1
NRC [45] 62.6 77.1 58.3 62.9 81.3 60.7 64.7 69.4 58.7 69.4 65.8 78.7 67.5
GKD [34] 61.4 77.4 60.3 69.6 81.4 63.2 68.3 68.4 59.5 71.5 65.2 77.6 68.7
AdaCon [1] 60.8 74.8 55.9 62.2 78.3 58.2 63.1 68.1 55.6 67.1 66.0 75.4 65.4
CoWA [17] 64.6 80.6 60.6 66.2 79.8 60.8 69.0 67.2 60.0 69.0 65.8 79.9 68.6
PLUE [23] 59.8 74.0 56.0 61.6 78.5 57.9 61.6 65.9 53.8 67.5 64.3 76.0 64.7
TPDS [38] 62.9 77.1 59.8 65.6 79.0 61.5 66.4 67.0 58.2 68.6 64.3 75.3 67.1
LCFD-C-RN [40] 75.4 88.2 72.0 75.8 88.3 72.1 76.1 75.6 71.2 77.6 75.9 88.2 78.0
LCFD-C-B32 [40] 77.2 88.0 75.2 78.8 88.2 75.8 79.1 77.8 74.9 79.9 77.4 88.0 80.0
DIFO-C-RN [41] 73.8 89.0 69.4 74.0 88.7 70.1 74.8 74.6 69.6 74.7 74.3 88.0 76.7
DIFO-C-B32 [41] 76.6 87.2 74.9 80.0 87.4 75.6 80.8 77.3 75.5 80.5 76.7 87.3 80.0
LLaVA-34B [25]* (w/ STS) - 84.4 91.0 83.7 85.5 91.0 83.7 85.5 84.4 83.7 85.5 84.4 91.0 86.1
InstBLIP-XXL [4]* (w/ STS) - 82.5 89.0 83.0 86.7 89.0 83.0 86.7 82.5 83.0 86.7 82.5 89.0 85.3
ShrGPT4V-13B [2]* (w/ STS) - 79.7 87.9 79.2 79.9 87.9 79.2 79.9 79.7 79.2 79.9 79.7 87.9 81.7
5.1 Main Results
Tables 2,  3, and  4 present our results for Office-Home, DomainNet, and VisDA, respectively. From top to bottom, we report domain adaptation methods that (1) use source data with CLIP-based techniques, (2) are source-free without multimodal or CLIP, and (3) employ CLIP-based multimodal approaches. Zero-shot MLLM (w/ STS)/CLIP performance is also included as a reference. RCL consistently achieves state-of-the-art performance across all datasets, with notable improvements: +6.4% on Office-Home, +9.4% on DomainNet, and +2.9% on VisDA-C. The prior best-performing methods, DIFO-C-B32 [41] and DIFO-C-RN [41], employ a ViT-B/32 CLIP encoder and ResNet CLIP backbone, respectively, using prompt learning. In contrast, RCL surpasses these methods while using only a ResNet backbone with guided curriculum training through MLLM pseudo-labels—without prompt learning or additional tuning. Similarly, PSAT-GDA [39], another competitive method, trains transformers for source guidance and domain alignment, whereas RCL relies solely on zero-shot MLLM inference and requires no additional tuning or training of large VLMs. Our self-refinement and curriculum learning processes outperform standard MLLM zero-shot performance by capturing valuable latent information beyond MLLMs. Additionally, the reliability of our pseudo-labels enables full-data training, further distinguishing RCL from prior approaches that rely on prompt tuning or handcrafted domain alignment. RCL leverages MLLM inference with STS and does not require customization, prompt learning, or heavy training of multimodal models.
Table 4: Accuracy (%) on VisDA-C dataset.
Source - 60.4 22.5 44.8 73.4 60.6 3.28 81.3 22.1 62.2 24.8 83.7 4.81 45.3
DAPL-RN [7] 97.8 83.1 88.8 77.9 97.4 91.5 94.2 79.7 88.6 89.3 92.5 62.0 86.9
PADCLIP-RN [15] 96.7 88.8 87.0 82.8 97.1 93.0 91.3 83.0 95.5 91.8 91.5 63.0 88.5
ADCLIP-RN [33] 98.1 83.6 91.2 76.6 98.1 93.4 96.0 81.4 86.4 91.5 92.1 64.2 87.7
SHOT [20] 95.0 87.4 80.9 57.6 93.9 94.1 79.4 80.4 90.9 89.8 85.8 57.5 82.7
NRC [45] 96.8 91.3 82.4 62.4 96.2 95.9 86.1 90.7 94.8 94.1 90.4 59.7 85.9
GKD [34] 95.3 87.6 81.7 58.1 93.9 94.0 80.0 80.0 91.2 91.0 86.9 56.1 83.0
AaD [47] 97.4 90.5 80.8 76.2 97.3 96.1 89.8 82.9 95.5 93.0 92.0 64.7 88.0
AdaCon [1] 97.0 84.7 84.0 77.3 96.7 93.8 91.9 84.8 94.3 93.1 94.1 49.7 86.8
CoWA [17] 96.2 89.7 83.9 73.8 96.4 97.4 89.3 86.8 94.6 92.1 88.7 53.8 86.9
SCLM [36] 97.1 90.7 85.6 62.0 97.3 94.6 81.8 84.3 93.6 92.8 88.0 55.9 85.3
ELR [48] 97.1 89.7 82.7 62.0 96.2 97.0 87.6 81.2 93.7 94.1 90.2 58.6 85.8
PLUE [23] 94.4 91.7 89.0 70.5 96.6 94.9 92.2 88.8 92.9 95.3 91.4 61.6 88.3
TPDS [38] 97.6 91.5 89.7 83.4 97.5 96.3 92.2 82.4 96.0 94.1 90.9 40.4 87.6
C-SFDA [13] 97.6 88.8 86.1 72.2 97.2 94.4 92.1 84.7 93.0 90.7 93.1 63.5 87.8
PSAT-GDA [39] 97.5 92.4 89.9 72.5 98.2 96.5 89.3 55.6 95.7 98.2 95.3 54.8 86.3
DIFO-C-RN [41] 97.7 87.6 90.5 83.6 96.7 95.8 94.8 74.1 92.4 93.8 92.9 65.5 88.8
DIFO-C-B32 [41] 97.5 89.0 90.8 83.5 97.8 97.3 93.2 83.5 95.2 96.8 93.7 65.9 90.3
LLaVA-34B [25]* (w/ STS) - 99.4 97.3 94.8 83.9 98.9 95.8 95.9 80.9 92.7 98.8 97.4 68.9 92.1
InstBLIP-XXL [4]*(w/ STS) - 99.2 89.6 82.0 69.8 97.9 91.0 97.5 84.3 73.6 99.3 96.7 60.0 86.7
ShrGPT4V-13B [2]*(w/ STS) - 99.2 94.7 90.8 87.9 98.3 92.1 97.3 68.0 96.3 95.6 96.8 68.2 90.4
5.2 Ablation Studies
This section covers ablation studies on RCL across its components, MLLMs used, backbones, and hyperparameters.
Table 5: Impact of RCL components (RKT, SMKE, MMR).
RCL Office-Home
[…]
82.8 73.3 89.3 88.1 83.3
88.5 80.9 95.1 92.5 89.3
87.7 80.2 93.3 92.0 88.3
89.3 82.3 95.3 92.9 90.0
5.2.1 Impact of RCL components.
Table 5 shows the results of evaluating individual RCL components. Using only RKT yields the lowest performance, as it relies on the most reliable MLLM pseudo-labels, which may lack class coverage and diversity (see Supplementary). Nonetheless, RKT provides sufficient initial supervision for identifying essential features. Applying SMKE after RKT outperforms MMR after RKT, as SMKE leverages less reliable pseudo-labels, improving robustness and expanding knowledge. In contrast, MMR directly after RKT performs worse, indicating the need for pseudo-label diversity before semi-supervised integration. Finally, MMR following SMKE consistently improves performance, allowing the model to learn from even the most unreliable labels in a semi-supervised manner, maximizing dataset utilization. Figure 8 visually compares the feature distributions of SOTA methods (DIFO) and RCL, showing RKT, SMKE, and MMR progressively refine target features.
Table 6: Performance of RCL with and without MLLMs on the Office-Home dataset. RCL (w/o MLLMs) uses three baseline methods (TPDS, LCFD-C-B32, and DIFO-C-B32).
[…]
TPDS 59.1 81.7 81.7 71.6 73.5
LCFD-C-B32 72.2 90.2 89.7 81.0 83.3
DIFO-C-B32 70.4 90.8 88.8 82.3 83.1
RCL (w/o MLLMs) 71.9 90.7 89.2 81.7 83.4
RCL (w/ MLLMs) 82.3 95.3 92.9 89.3 90.0
5.2.2 Synergy between RCL and MLLMs.
Table 6 shows that without MLLMs, RCL provides only a marginal 0.1% gain over the best existing SFDA method (LCFD-C-B32), indicating limited improvement when applied to standard adaptation approaches (TPDS, LCFD-C-B32, DIFO-C-B32). In contrast, integrating MLLMs (LLaVA-34B, ShareGPT4V-13B, InstBLIP-XXL) into RCL yields an 2.8% improvement over the best MLLM model (LLaVa-34B) indicated in Table 2 and a 6.4% boost over RCL without MLLMs. Unlike ImageNet-pretrained models, MLLMs are trained on broad multimodal corpora, providing valuable auxiliary knowledge that enhances pseudo-label quality and adaptation effectiveness. These results confirm that traditional SFDA methods lack the generalization capacity of MLLMs, while RCL effectively leverages their complementary knowledge for improved adaptation.
Figure 5: RCL’s sensitivity and robustness against MLLMs with weaker capability and MLLM ensemble.
5.2.3 Sensitivity to the capability of MLLMs.
Figure 5 compares two settings: (1) weaker MLLMs with lower zero-shot performance and (2) the strongest MLLMs with the highest ensemble accuracy. Labels are determined by majority vote, with ties assigned randomly. RCL consistently surpasses individual MLLMs and their ensemble, with the gap most pronounced in Setting 1 (4.3% increase), where weaker MLLMs struggle. Even in Setting 2, where MLLMs perform well, RCL achieves superior results over their ensemble (1.4% increase). These results confirm RCL’s ability to consistently enhance MLLM performance, especially when MLLMs have low individual performance.
5.2.4 Number of MLLMs.
Figure 6 shows the impact of MLLM count on RCL performance, comparing using a single MLLM to four. We use the strongest models individually (ShareGPT-13B, InstructBLIP-XXL, LLaVA-34B), combining them for three, and adding BLIP2-XXL for four. A single MLLM yields the lowest accuracy (see Figure 7) since RKT-only learning is constrained by the MLLM-STS zero-shot upper bound. In contrast, RCL surpasses each MLLM by integrating insights and adapting knowledge. Comparing three vs. four MLLMs, we find no significant difference, indicating RCL efficiently leverages multiple MLLMs without requiring a large ensemble.
Table 7: Ablation study on the choice of backbone (BB).
Method BB Office-Home
[…]
DIFO-C-RN RN50 79.3 63.1 87.7 87.5 79.4
DIFO-C-B32 RN50 82.3 70.4 90.8 88.3 83.1
RCL (Ours) RN18 89.1 81.5 95.1 92.6 89.6
RCL (Ours) RN50 89.3 82.3 95.3 92.9 90.0
[…]
5.2.5 Knowledge transfer to a smaller backbone.
We investigate transferring to a smaller backbone for the scalability of SFDA. As shown in Table 7, RCL achieves similar performance with ResNet18 as with ResNet50, maintaining a +6.4% accuracy advantage over state-of-the-art methods. With twice the speed (2 GFLOPS v.s. 4 GFLOPS), RCL effectively distills pre-trained knowledge to a smaller backbone, making it ideal for large-scale inference across diverse tasks. This demonstrates RCL’s capability to efficiently leverage MLLM knowledge and adapt it to a lightweight model suitable for real-world deployment.
Table 8: Average Latency (ms / sample) comparison for inference. RCL uses RN50 backbone.
Model Avg. Latency $\downarrow$
LLaVA-34B (w/ STS) $\sim$2850
ShrGPT4v-13B (w/ STS) $\sim$1890
InsBLIP-XXL (w/ STS) $\sim$2740
RCL $\sim$5
5.2.6 Latency of deployed model.
Table 8 illustrates the inference latency comparison of MLLMs and our proposed RCL. While LLaVA-34B, InstructBLIP-XXL, and ShareGPT4v-13B require latency over 1800ms per sample, RCL achieves a significantly faster inference speed of  5ms per sample. When combining all three MLLMs for ensembling, over 7000ms latency is required. This over 378x improvement in computational efficiency validates our approach of distilling MLLM knowledge into a compact model rather than relying on direct MLLM inference, enabling practical deployment while achieving even better performance. All SFDA methods using ResNet-50 have comparable inference latency. RCL adds complexity only during initial pseudo-label generation with MLLMs but maintains efficient training and inference post-deployment while achieving superior performance. In contrast, DIFO [41] incorporates additional models (e.g., CLIP-RN, CLIP-ViT) in every training iteration, increasing computational overhead throughout training.
Figure 9: The effect of $\tau$ in SMKE.
5.2.7 Sensitivity of hyper-parameters.
Figure 9 illustrates the effect of the confidence threshold $\tau$ in SMKE. It shows how varying $\tau$ values influence model performance by balancing self-correction with MLLM guidance. Higher $\tau$ values increase reliance on high-confidence pseudo-labels from the target model, while lower values depend more on MLLM-generated pseudo-labels. The figure highlights the optimal $\tau$ that enhances adaptation performance by effectively leveraging both reliable and less reliable pseudo-labels. For the MMR, Table 9 shows that $\tau$ = 0.95 provided a slight edge in accuracy, ensuring the model effectively utilizes its predictions while still considering MLLM guidance. In Table 10, we evaluate $\lambda$ values from 0 to 2.0. The results show that $\lambda$ = 0.5 consistently yielded the highest average accuracy across multiple domain adaptation tasks, suggesting it strikes a good balance between supervised learning and regularization.
Table 9: The effect of confidence threshold $\tau$ in MMR.
$\tau$ 1.0 0.95 0.85 0.75 0.65
[…]
Table 10: Effect of different $\lambda$ values in MMR.
[…]
0 82.0 95.3 92.8 89.1 89.8
0.5 82.3 95.3 92.9 89.3 90.0
1 82.3 95.3 92.8 89.2 89.9
1.5 82.2 95.2 92.8 89.2 89.9
2 82.3 95.2 92.8 89.1 89.9
5.2.8 MMR Complexity and Effectiveness.
MMR demonstrates enhanced effectiveness for challenging adaptation tasks, with performance gains correlating to the proportion of unreliable samples. For example, the ($\rightarrow$ Clipart) achieves the most significant improvement of +1.4% (80.9% to 82.3% in Table 5), despite having the highest proportion of unreliable (R=0) samples (see Fig. 10). These results indicate MMR’s particular utility when dealing with large portions of unreliable labels, which aligns with real-world scenarios where MLLMs possess divergent knowledge spaces. MMR enables comprehensive data utilization through its ability to learn from all samples, rather than being limited to only high-confidence instances. Our lightweight inference model ensures deployment efficiency, achieving a practical latency of 5ms per sample.
Regarding the effectiveness of MMR components, we have conducted detailed ablation studies (Tables 5 and  10). These results demonstrate that both multi-hot masking and consistency loss contribute to the performance. Note that the MMR cannot be applied without Multi-hot Masking since the unreliable samples cannot be used by traditional pseudo-labeling methods.
Figure 10: Per-class counts of MLLM pseudolabels on Office-Home.
6 Conclusion
We introduce a novel approach for adapting foundation knowledge MLLMs to significantly enhance SFDA, transforming their zero-shot capabilities into a structured adaptation process. Since MLLMs are primarily designed for text generation and may produce inconsistent outputs, we first introduce STS to align their zero-shot predictions with target classes, enabling their use as reliable pseudo-labelers in SFDA. Building on this, we propose RCL, which systematically leverages multiple MLLMs to refine pseudo-labels and improve adaptation. RCL employs curriculum learning to dynamically segment pseudo-labels by reliability, progressively refining knowledge and adaptation. This enables a self-correcting learning process that distills MLLM knowledge into a lightweight model, making it more practical for real-world deployment. While our focus is on using MLLMs for SFDA enhancement, this work serves as a broader demonstration of how foundation MLLM knowledge can enhance vision tasks and beyond.
Limitation. Although RCL’s generalizability may still be affected by biases inherent in pre-trained MLLMs, our consensus-based approach attempts to mitigate this by aggregating knowledge from multiple MLLMs, reducing dependence on any single model’s bias.

### v3

@@ 9 Multi-hot Masking Refinement Details
10 MLLM Pseudo-labeling Details
In this section, we provide details on prompt templates (Sec. 10.1), and discussions on pseudolabel distribution (Sec. 10.3) and prompt sensitivity (Sec. 10.2).
10.1 Prompt Templates
To facilitate pseudo-labeling, we design specific prompt templates for different MLLMs:
LLaVA models: "Question: What is the closest name from this list to describe the object in the image? Return the name only. <class names>"
ShareGPT4V models: "Question: What is the closest name from this list to describe the object in the image? List: <class names> Return the closest name from the list only. Use *exact* names from the list only. Answer:"
InstructBLIP models: "Question: What is the closest name from this list to describe the object in the image? <class names>. Use the closest name from the list only. Pick the answer from the list only. Answer:"
For the STS calculations, we utilize all-MiniLM-L6-v2 to generate vector representations of both the outputs from the MLLMs and the text options [reimers-2020-multilingual-sentence-bert].
10.2 Prompt Sensitivity
To evaluate robustness of our STS strategy, we tested LLaVA-34B on OfficeHome under multiple prompt templates. Beyond the default template, we considered (i) a variant with added domain information, (ii) a naive template phrased as “What is this image from given list below {class_list}?”, (iii) a paraphrased class list, and (iv) a list with injected distractor labels. For paraphrased class names, we manually curated synonyms or simplified forms of Office-Home categories (e.g., “alarm clock” → “clock”, “computer monitor” → “monitor”, “filing cabinet” → “cabinet”, “push pin” → “thumbtack”). While MLLM outputs differ since they pick from modified class lists, STS reliably maps predictions back to the original label set, showing consistency under lexical variation. For distractor labels, we manually injected unrelated category names sampled from DomainNet and VisDA into the candidate list, simulating noisier prompts where irrelevant options appear alongside correct classes. These variations stress-test both STS alignment and consensus-driven reliability, showing that RCL remains stable despite prompt perturbations.
Results are shown in Table 17. Across all four Office-Home transfers, performance varies by less than 2% absolute between prompt templates. While domain information slightly reduces accuracy, and paraphrases/distractors introduce minor drops, the overall averages remain stable.
Table 17: Prompt sensitivity analysis on Office-Home using STS. Zero-shot+STS accuracy with different prompt templates for LLaVA-34B show small fluctuations.
Prompt Template Ar→Cl Ar→Pr Ar→Rw Avg.
Naive prompt 76.80 92.10 87.45 85.71
Default prompt (D) 78.35 93.78 89.58 87.19
D + Domain info 77.54 93.08 88.73 86.25
D + Paraphrased classes 76.20 92.45 87.90 85.20
D + Distractor labels 76.90 92.70 88.25 85.79
10.3 Pseudo-label Distributions
Fig. 10 demonstrates the distribution of sample reliability across different classes within the Office-Home dataset. We observe that for various classes across domains, there are few samples with R=1 (high reliability). This emphasizes the importance of utilizing less reliable data, a fundamental aspect of our approach, which is integrated into our SMKE and MMR techniques. To highlight our point, if we observe Table 3, it is evident that using only R=1 data (applying only RKT) results in notably lower accuracies for the Clipart domain (while adaptations to other domains achieve over 80%, Clipart is lowest at 73.3%). As Fig. 10 indicates, the Clipart domain has a higher proportion of pseudo-labels where 0<R<1. By employing SMKE and MMR, we manage to enhance performance in adapting to the Clipart domain by +9% (to 82.3%). We also provide the distributions for $D_{R}$, $D_{LR}$, and $D_{UR}$ across various teacher ensembles and respective RCL accuracy in Table 14.
Figure 10: Per-class counts of pseudo-labels on Office-Home using LLaVA-34B, InstructBLIP-XXL, and ShareGPT4v-13B.
11 Extended Related Work on SFDA and DA
SFDA Methods. Early SFDA methods adapted ImageNet-pretrained models using pseudo-label refinement, such as SHOT [SHOT_2021], NRC [NRC_2022], and G-SFDA [yang2021generalized]. These works emphasize clustering, entropy minimization, or distribution alignment under the no-source setting. More recent SFDA studies refine pseudo-labeling with neighborhood consistency [tang2021nearest, tang2022semantic] or target-structure guidance [chenself, litrico2023guiding]. TPDS [lai2024empowering] introduced target-prototype distribution smoothing for improved calibration but still relies on strong pre-trained features (e.g., CLIP). DIFO [tang2023sourceMLLM], the current SOTA, adapts CLIP through prompt learning and distillation, progressively aligning CLIP with target data before transferring to a task-specific model. While effective, these works typically exploit a single teacher or finetune a large VLM, making them less scalable when source data is unavailable or when deploying lightweight students is required.
Prompt-based DA with CLIP. Although not strictly SFDA, prompt-based DA methods highlight another direction. Chen et al. [chen2023multi] proposed multi-prompt alignment for multi-source DA; Phan et al. [hoang2024enhancing] improved cross-domain transfer by aligning prompt gradients; Ge et al. [ge2023domain] developed prompt learning for domain adaptation; and Vuong et al. [vuong2025preserving] preserved clusters in prompt learning for UDA. These approaches require source data or the ability to tune CLIP prompts, unlike SFDA where such access is not allowed. We include them here to show that our reliability-driven framework complements prompt-based DA by demonstrating how frozen teachers (MLLMs or CLIP) can be exploited without modification.
Positioning of Our Work. Our RCL framework sits at the intersection of these lines: (1) like SFDA methods, it respects the no-source constraint, but unlike prior work, it incorporates multiple teachers; (2) like CLIP-based DA methods, it leverages large pretrained multimodal models, but without prompt tuning or finetuning; and (3) unlike DIFO or TPDS, it distills supervision into compact students by explicitly modeling teacher reliability and progressively incorporating all target data through a curriculum. This distinction is critical: RCL is not limited to a single CLIP teacher or large MLLMs, but provides a general strategy for multi-teacher SFDA under noisy pseudo-labels.
12 Extended Discussions
12.1 Why use MLLMs?
MLLMs as Teachers. A key motivation for adopting MLLMs is their role as foundation models that pair strong visual encoders with instruction-tuned LLMs, yielding broader semantics and instruction-following ability than contrastive VLMs like CLIP. Large-scale visual instruction tuning has produced general-purpose MLLMs with strong zero-shot performance across diverse image understanding tasks, as shown in multi-discipline benchmarks [meng2024mmiu, wang2024muirbench, li2024seed, chen2024mllm, yue2024mmmu]. Beyond general benchmarks, domain-specific MLLMs such as LLaVA-Med [li2023llava-med] and biomedical/industrial evaluations like MicroVQA [lozano2024micro], and others [lai2025can, sepehri2024mediconfusion], show their practical availability where task-specific models are scarce. Importantly, recent studies show that MLLMs have now caught up with and in many cases surpassed CLIP on standard image classification benchmarks [liu2024revisiting], highlighting their promise as future-proof teachers for SFDA. RCL provides a recipe to distill such broad supervision into compact, task-specific students without prompt tuning, teacher training, multiple inferences per sample, or source data access. In our pipeline, noisy MLLM predictions or hallucinations are naturally handled as unreliable labels, mitigated through reliability-based multi-teacher distillation.
Robustness to Teacher Variants. At the same time, RCL is not contingent on large MLLMs. Our framework is fundamentally teacher-agnostic, focusing on multi-teacher curriculum learning under noisy pseudo-labels. Stress tests confirm this: replacing an MLLM with CLIP-RN50 reduces Office-Home accuracy by only 1.3% (Supp. Table 10), while ensembles of weaker or smaller teachers (72.9% average zero-shot accuracy) still achieve 81.1% with RCL, surpassing strong CLIP-only SFDA baselines. Moreover, with modest student backbones (RN18/50), RCL outperforms ViT-based CLIP methods (Table 4, main), showing that the gains stem from reliability-driven distillation rather than teacher scale. Thus, MLLMs serve as a strong upper-bound case that demonstrates the potential of foundation models for SFDA, while RCL remains robust and effective with CLIP or heterogeneous teacher ensembles. We summarize key differences in using MLLM vs. CLIP/VLM models as techers in Table 18.
12.2 Practicality of RCL.
Although MLLMs provide strong zero-shot classification, their direct use for classification is computationally expensive and unsuitable for large-scale or real-time deployment. RCL addresses this by requiring only a single inference pass of MLLMs (or other teachers) over the target dataset. These pseudo-labels are cached and used throughout training, meaning the heavy models need not be invoked again. In contrast, CLIP-based prompt learning often requires repeated forward passes per image across multiple prompts, and direct MLLM inference is prohibitive for deployment. By distilling cached pseudo-labels into a lightweight target model, RCL produces compact students that retain the benefits of foundation knowledge while being orders of magnitude cheaper to train and deploy. This design enables real-world SFDA usage: practitioners can harness the strengths of foundation models without maintaining them in production, instead deploying efficient students adapted with RCL.
Table 18: Motivation and comparisons for MLLM vs. CLIP/VLM as teacher models.
Category MLLM Teachers CLIP / VLM Teachers
Efficacy Richer semantics, reasoning, instruction-following. Zero-shot MLLMs w/ STS already surpass CLIP baselines (Tab. 1,2). RCL w/ MLLMs improves +6.4% over CLIP-only RCL (Tab. 4). Strong contrastive features, but limited to alignment. Require prompt tuning / finetuning for strong SFDA (e.g., DIFO). RCL with CLIP substitution still competitive (Supp. Tab. 10).
Computation High per-query latency (1800–2800ms, Supp. Tab. 9). In RCL, used once for caching pseudo-labels; no repeated inference. Lower latency per query, but prompt-learning methods require repeated inference per epoch, increasing total cost.
Deployment Not suitable for on-device deployment due to size. Distillation via RCL yields compact students ($\sim$5ms/sample). More deployable, but adaptation pipelines (e.g., DIFO) still depend on CLIP during training, raising training-time overhead.
Robustness Aggregating diverse MLLMs mitigates hallucination; reliability scoring filters noise. Gains persist even with weaker MLLMs (Supp. Fig. 6, Tab. 10). Ensembles can cover variance, but raw ensembling underperforms RCL distillation (Supp. Fig. 8).
Generalizability Encodes broad, multi-domain knowledge (general + domain-specific MLLMs). Strong upper bound for SFDA and beyond. Strong on vision tasks but limited cross-domain reasoning. Less effective in medical/industrial applications where MLLM knowledge is valuable.
12.3 RCL as a General Multi-Teacher SFDA Framework
While we highlight MLLMs in the main paper, RCL can be seen as a general multi-teacher curriculum distillation framework for SFDA. The core components (RKT, SMKE, and MMR) are teacher-agnostic, working on pseudo-label reliability rather than specific architectures. Our proposed STS is what makes MLLMs viable teachers by converting free-form text outputs into closed-set predictions; for models like CLIP or ViTs, this conversion is not required and the same RCL can be applied.
We validate this generality through multiple experiments:
  1. 1.
    CLIP substitution in Table 13. Replacing an MLLM with CLIP-RN50 reduces accuracy by only 1.3%.
  2. 2.
    Weaker/smaller teachers in Table 13. Even ensembles with an average of 72.9% zero-shot accuracy achieve 81.1% with RCL, surpassing strong CLIP-based SFDA methods.
  3. 3.
    Comparison to ViT-based baselines (TPDS [tang2023source], LCFD [tang2024unified], DIFO [tang2023sourceMLLM]) in Table 4 (main paper). RCL-trained ResNet student outperforms ViT-based CLIP prompt-learning methods, highlighting that the framework (not teacher scale) is the main driver.
Thus, while MLLMs provide richer semantics and broad applicability (e.g., medical, industrial domains where task-specific teachers are unavailable), RCL is not confined to MLLMs. It is a robust, scalable framework that can integrate diverse teacher families, making it broadly useful for SFDA across settings. Future work can assess its viability and adoption across different tasks beyond SFDA.